OpenStudy (anonymous):

The following table shows the length and width of two rectangles: Length Width Rectangle A x + 8 x − 1 Rectangle B 4x + 5 3x − 2 Which expression is the result of the perimeter of rectangle B minus the perimeter of rectangle A? 10x − 12 18x + 32 10x − 8 18x + 6

OpenStudy (anonymous):

can someone help me on this?? i dont understand it

OpenStudy (jtvatsim):

Yikes! A little scary, huh? Not to worry though, it's actually pretty easy. Let's take it one step at a time.

OpenStudy (anonymous):

ok great

OpenStudy (jtvatsim):

The question sounds very confusing, "the perimeter of rectangle B minus the perimeter of rectangle A"... uh.. yeah. Let's just ignore that for now and think about the perimeters separately: Perimeter of rectangle A is (we will find out) and Perimeter of rectangle B is (again, we will find that out).

OpenStudy (anonymous):

okay

OpenStudy (jtvatsim):

So, what do we know about rectangle A? The table says that A's length is x + 8 (some number) and the width is x - 1 (some other number). Let's draw a picture so we can see what this even means.

OpenStudy (jtvatsim):



OpenStudy (jtvatsim):

Alright, does that picture make sense so far?

OpenStudy (anonymous):

yes, it does

OpenStudy (jtvatsim):

OK, perfect! We are almost done with the hardest part. Do you remember how to find the perimeter?

OpenStudy (anonymous):

you time it by each side ??

OpenStudy (anonymous):

times*

OpenStudy (jtvatsim):

Close, that's the area (they often get confused). :)

OpenStudy (jtvatsim):

It's actually easier... we just add all four sides together and measure the outside of the rectangle.

OpenStudy (anonymous):

ooh okay

OpenStudy (jtvatsim):

(x+8) + (x-1) + (x+8) + (x-1) = 4x + 16 - 2 = 4x + 14. There that's better.

OpenStudy (jtvatsim):

So that is the perimeter of A. Any questions on how we did that part?

OpenStudy (anonymous):

nope

OpenStudy (jtvatsim):

OK. Now, we just do exactly the same thing for rectangle B.

OpenStudy (jtvatsim):

We have different numbers: B width = 3x - 2 B length = 4x + 5

OpenStudy (jtvatsim):



OpenStudy (anonymous):

awesome

OpenStudy (jtvatsim):

And, then we just add all four sides... it's a little messy, but try it and see what you get. I'll check your answer. :)

OpenStudy (anonymous):

ok give me a second

OpenStudy (jtvatsim):

sure thing. take your time. :)

OpenStudy (anonymous):

4x+5+3x-2+4x+5+3x-2= 14x+6

OpenStudy (jtvatsim):

Excellent! Good job!

OpenStudy (jtvatsim):

Alright, so now we know that: Perimeter of A = 4x + 14 Perimeter of B = 14x + 6.

OpenStudy (anonymous):

so subtract 14x+6- 4x+14

OpenStudy (anonymous):

?

OpenStudy (jtvatsim):

Good, but be careful with your parentheses. We want the whole perimeter of A to be subtracted so we need (14x + 6) - (4x + 14) in order to get the right answer.

OpenStudy (anonymous):

so its 10x-8 ??

OpenStudy (jtvatsim):

Exactly!

OpenStudy (anonymous):

YAYYY THANKYOU SO MUCH :))

OpenStudy (jtvatsim):

You're welcome! Good job!

OpenStudy (anonymous):

:)